# ARC: Argument Representation and Coverage Analysis for Zero-Shot Long Document Summarization with Instruction Following LLMs

### v1

@@ Appendix G Correlation with source argument positions across full ARC scores
Appendix H Bias analysis for argument coverage without argument role normalization
@@
While our normalization in computing corrects for frequency skew, it may understate coverage for dominant argument roles in the source with inherently high raw scores. Therefore, we also examine role-specific reporting bias by directly computing to offer a complete picture of the role bias analysis. As shown in Figure 10, results on CANLII confirm prior findings: LLMs tend to prioritize covering conclusion arguments over issue and reason roles, both in the controlled and non-controlled length settings. For DRI, removing frequency normalization reveals that ARC scores for roles such as background claim, own claim, and data are comparable to each other. This suggests that without normalization, the higher coverage of own claim may stem from its over-representation in the source documents, rather than a model-level preference.

### v2

@@ Appendix H Source sentences identification
Appendix I Bias analysis for argument coverage without argument role normalization
@@
While normalization in computing corrects for frequency skew, it may also understate coverage for dominant roles with inherently high raw scores. To provide a fuller picture, we additionally compute role-specific bias directly as . As shown in Figure 7, results on CANLII100 confirm prior findings: LLMs disproportionately prioritize conclusions over issues and reasons, both under length-controlled and non-controlled settings. For DRI, removing frequency normalization explains partially the higher scores reported in Section 5.4.
Appendix J Precision–recall trade-offs in ARC
ARC is designed to prioritize recall, aligning with prior coverage-based factuality metrics (e.g., FactScore) that emphasize capturing as many salient facts as possible in the summary. Nevertheless, we believe that incorporating precision-based analyses provides a more complete view of different models behavior. To this end, we conducted an additional study on summaries from the CANLII dataset, evenly sampled across summary length buckets. Using the same atomic fact decomposition and entailment filtering setup as in the main experiments, we compute: Precision: proportion of covered salient facts divided by the number of facts generated in the summary. Recall: proportion of reference salient facts covered by the summary. F1: harmonic mean of precision and recall. Both atomic fact decomposition and evaluation were conducted using GPT4-o.
Table 18 reports precision, recall, and F1 scores across models.
| Model | Precision | Recall | F1 |
| Qwen-2.5-14B | 0.8766 | 0.6418 | 0.7149 |
| Qwen-2.5-7B | 0.6337 | 0.6374 | 0.5737 |
| Qwen-2.5-3B | 0.5027 | 0.5874 | 0.5305 |
| Qwen-2.5-1.5B | 0.7733 | 0.5608 | 0.6227 |
| Mistral-8B | 0.7695 | 0.5188 | 0.5724 |
| LLaMA-3.1-8B | 0.7734 | 0.6174 | 0.6473 |
| LLaMA-3.2-3B | 0.5622 | 0.3983 | 0.4354 |
| LLaMA-3.2-1B | 0.4618 | 0.4069 | 0.4080 |
To further understand how evaluation choice affects model comparison, Table 19 shows model rankings induced by each metric.
| Model | Precision Rank | Recall Rank | F1 Rank |
| Qwen-2.5-14B | 1 | 1 | 1 |
| Qwen-2.5-7B | 5 | 3 | 5 |
| Qwen-2.5-3B | 6 | 5 | 4 |
| Qwen-2.5-1.5B | 3 | 4 | 3 |
| Mistral-8B | 4 | 6 | 6 |
| LLaMA-3.1-8B | 2 | 2 | 2 |
| LLaMA-3.2-3B | 7 | 8 | 7 |
| LLaMA-3.2-1B | 8 | 7 | 8 |
Finally, Table 20 reports rank correlation coefficients between rankings induced by different metrics.
| Metric Pair | Kendall | Pearson |
| Precision vs. Recall | 0.50 | 0.62 |
| Precision vs. F1 | 0.86 | 0.95 |
| Recall vs. F1 | 0.64 | 0.81 |
Overall, while precision introduces moderate reordering among models, the overall relative ranking of models remains largely stable. The positive correlations between recall and precision indicate that ARC’s recall-oriented design captures a general performance trend that largely persists even under precision-aware evaluation.